The half-life of radium is 1600 years. The fraction of a sample of radium that would remain after 6400 years


1
Expert's answer
2021-06-29T03:54:03-0400

The fraction of a sample that would remain after the time "t" is given by the law of radioactive decay:


"\dfrac{N}{N_0} = 2^{-\dfrac{t}{T}}"

Substituting "t = 6400years" and the half-life "T = 1600years", obtain:


"\dfrac{N}{N_0} = 2^{-\dfrac{6400}{1600}} = 0.0625"

Answer. 0.0625.